OpenStudy (anonymous):

1/1-3+1/1+3=10/x^2-9

OpenStudy (callisto):

do you mean (1/1-3) +(1/1+3) =10/(x^2-9) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

could you help me solve ?

OpenStudy (callisto):

(1/1-3) +(1/1+3) =10/(x^2-9) (-1/2) + 1/4 =10/(x^2-9) -1/4 = 10/(x^2-9) 9-x^2 = 40 x^2 = -31 ...... \[x=\pm\sqrt{31} i\]